OpenStudy (anonymous):

SOMEBODY PLZ HELP ME ON THIS?? Bill Palow invested 30,000 for one year in two separate accounts paying 3% and 4.1% annual simple interest. If Bill earned a total of 1091.73 from the two investments, how much was invested in each account?

OpenStudy (anonymous):

I=PRn let R1=0.03 let R2=0.041 Invested amount = 30000 = a+b I=1091.73=1(0.03a + 0.041b) a = 30000-b Sub 1091.73=0.03(30000-b)+0.041b

OpenStudy (anonymous):

Find b and then sub to find a

OpenStudy (anonymous):

I know I have terrible setting out

OpenStudy (anonymous):

THE N IN THE EQUATION STANDS FOR NUMBER OF YEARS RIGHT?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

SO THE PRINCIPAL EQUALS 3OOOO?

OpenStudy (anonymous):

P=a+b

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

SO THEN I EQUALS 30000

OpenStudy (anonymous):

P= invested amount I = interest, which is amount that is earned n = number of years R = rate of interest per annum So yes P =30000

OpenStudy (anonymous):

THNK YOU